OpenStudy (anonymous):

Factor by grouping. 3x2 + 7x – 6 A. (3x – 2)(x – 3) B. (3x – 2)(x + 3) C. (x + 3)(3x + 2) D. (3x + 2)(x – 3)

OpenStudy (anonymous):

I've eliminated A. and C.... then I got stuck

hartnn (hartnn):

to do factor by grouping of the standard quadratic expression,\( ax^2+bx+c\), we need 2 numbers whose sum is 'b' and product = ac for your Question, can you find 2 numbers whose sum =7 and product = -6*3=-18 ?

OpenStudy (ghostfire):

B. (3x-2)(x+3) =\[3x ^{2}+9x-2x-6\] D. (3x+2)(x-3) =\[3x ^{2}-9x+2x-6\] so which one would give you +7x ?

OpenStudy (anonymous):

I'm sooooo confused >.<

hartnn (hartnn):

whats 9x-2x =... ?

OpenStudy (anonymous):

7x

OpenStudy (ghostfire):

so B is the answer, because answer D would give you -7x

OpenStudy (ghostfire):

Since this is a multiple choice question, if you don't know how to factor to get the answer. Do it backward by multiply the answer out and see which one match your question.

OpenStudy (anonymous):

Oooh, I see..thank you.